OpenStudy (anonymous):

Compute \frac{3}{4}\cdot \frac{8}{9}\cdot \frac{15}{16}\cdot \frac{24}{25}\cdot \frac{35}{36}\cdot \frac{48}{49}\cdot \frac{63}{64}\c.... Express your answer in the simplest way possible.

OpenStudy (agreene):

\[\frac{3}{4}\cdot \frac{8}{9}\cdot \frac{15}{16}\cdot \frac{24}{25}\cdot \frac{35}{36}\cdot \frac{48}{49}\cdot \frac{63}{64}\]

OpenStudy (anonymous):

Latex doesn't work!!

OpenStudy (agreene):

fix'd that for you

OpenStudy (anonymous):

It's \[\prod_{i=1}^\infty \dfrac{i^2+2i}{i^2+2i+1} \]

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

let check: i =1, then, numerator = 3, denominator = 4 i =2, then , numerator = 8, denominator =9 i =3, then numerator = 15, denominator =16 and so on. the \(\prod\) stands for multiplication, right? , then it works well. hihihi

OpenStudy (mathmale):

Oooops! That's a clever and sophisticated approach. Another way (and a good check) would be to cancel everywhere possible in this fraction. For example, that 8 in the numerator, divided into that 16 in the denom., results in a 2 in the denom. And so on.